OpenStudy (katielong):

integrate, using the substitution u=e^2x +4, {2/(e^2x +4)

OpenStudy (katielong):

\[\int\limits_{.}^{.} \frac{ 2 }{ e ^{2x} +4} dx \]

OpenStudy (katielong):

\[u=e ^{2x}+4\]

OpenStudy (katielong):

substitution of u gives... \[\int\limits_{.}^{.}\frac{ 2 }{ u } dx\] and \[dx=\frac{ du }{ 2e ^{2x} }\]

OpenStudy (katielong):

therefore \[\int\limits_{.}^{.}\frac{ 2 }{ u } \frac{ du }{ 2e ^{2x} }\]

OpenStudy (katielong):

the twos cancel so...

OpenStudy (katielong):

\[\int\limits_{.}^{.} \frac{ 1 }{ ue ^{2x} } du\]

OpenStudy (katielong):

...but then I'm stuck, can you integrate ergardless of the fact there's still an x within th equation or do you have to substitute x for u?

OpenStudy (irishboy123):

e^2x = u - 4

OpenStudy (irishboy123):

IOW, you must switch it all to u...

OpenStudy (katielong):

oh right, I see

OpenStudy (katielong):

so you integrate: \[\int\limits_{.}^{.} \frac{ 1 }{ u(u-4) } du\] which is equal to~: \[\int\limits_{.}^{.}\frac{ 1 }{ u ^{2}-4u }\]

zepdrix (zepdrix):

Oh I was going to correct a mistake, but it looks like you've got it set up correctly! Hehe. So yah. Partial Fraction Decomposition from there I guess.

OpenStudy (katielong):

partial what what??

zepdrix (zepdrix):

\[\Large\rm \frac{1}{u(u-4)}=\frac{A}{u}+\frac{B}{u-4}\]Never learned this? :o

OpenStudy (katielong):

oh yeah i recognise that, so would A be (u-4) and B (u)?

zepdrix (zepdrix):

No no, A and B end up being just numbers.

zepdrix (zepdrix):

To figure out A and B, we'll multiply both sides through by that denominator on the left. Giving us,\[\Large\rm 1=A(u-4)+Bu\]Ok with that step? :o

OpenStudy (katielong):

sorry yes that right

OpenStudy (katielong):

A= -1/4 and B = 5/4?

zepdrix (zepdrix):

Hmm I think B=1/4... checking..

OpenStudy (katielong):

yeah sorry it is, i just did that quickly

zepdrix (zepdrix):

\[\Large\rm \int\limits \frac{1}{u(u-4)}du=\int\limits \frac{A}{u}+\frac{B}{u-4}~du\] \[\Large\rm =\int\limits \frac{-1/4}{u}+\frac{1/4}{u-4}~du\]So we get something like that, yah? :o

OpenStudy (katielong):

yep

OpenStudy (katielong):

\[\int\limits_{.}^{.} \frac{ -4 }{ u } + \frac{ 4 }{ u-4 }\]

zepdrix (zepdrix):

No no no silly :3 We only flip the fraction when it's in the bottom.

zepdrix (zepdrix):

\[\Large\rm =\int\limits\limits \frac{\color{orangered}{-1/4}}{u}+\frac{\color{orangered}{1/4}}{u-4}~du\]These are in the top :o

OpenStudy (katielong):

okayyy

zepdrix (zepdrix):

You could do this I suppose, if you want it to be easier to read, factor a 1/4 out of each term,\[\Large\rm =\frac{1}{4}\int\limits -\frac{1}{u}+\frac{1}{u-4}~du\]

OpenStudy (katielong):

ah okayy yeah

zepdrix (zepdrix):

Remember how to integrate those two guys? :D

OpenStudy (katielong):

1/4 [ -lnu + ln Iu-4I ] ?

zepdrix (zepdrix):

Mmm looks good. Now to undo our sub :)

OpenStudy (katielong):

-lnu/4 + (lnIu-4I)/4 ?

zepdrix (zepdrix):

We COULD distribute the 1/4 back in, but let's leave him out front for now maybe.. I think he's just gonna mess things up for us XD

OpenStudy (katielong):

okayy, sorry i misunderstood you before haha, so substitute u for e^2x +4 ?

zepdrix (zepdrix):

\[\Large\rm \frac{1}{4}\left[\ln(\color{orangered}{u-4})-\ln(\color{orangered}{u})\right]\]Yah :) gotta do somethign with that orange stuff.

OpenStudy (katielong):

u-4 = e^2x u = e^2x +4

zepdrix (zepdrix):

\[\Large\rm \frac{1}{4}\left[\ln(\color{orangered}{e^{2x}})-\ln(\color{orangered}{e^{2x}+4})\right]\]Sounds good so far :D

OpenStudy (katielong):

1/4 [ 2x - ln(e^2x +4 ]

zepdrix (zepdrix):

The first one simplifies? Ahh nice :D

zepdrix (zepdrix):

You can throw the 1/4 in now if you want I suppose :P But whatever, that's a fine looking answer. Make sure to throw a +c somewhere in the mix since it's an `indefinite` integral.

zepdrix (zepdrix):

Yayyy good job katie \c:/

OpenStudy (katielong):

okay thank youu, I'm soo bad at integration :( everything else I'm absolutely fine at but I missed a few lessons of integration and I'm just awful at it :L

zepdrix (zepdrix):

There are so many little tricks and tools for integration :O Takes a while to get comfortable with them. Do tons and tons of practice problems to get a feel for what approach you want to take.

zepdrix (zepdrix):

Like whenever you see a square in a denominator, or under a root, a lightbulb should go off in your head saying "oh probably a trig sub"

zepdrix (zepdrix):

A square with addition/subtraction*

OpenStudy (katielong):

yeah I certainly need too! The text book I have is awful though!

OpenStudy (katielong):

Are there any good websites to help me or anything?

zepdrix (zepdrix):

Hmm there's probably a lot of cool stuff out there :d I found this website to be really helpful when going through Differential Calculus: https://www.math.ucdavis.edu/~kouba/ProblemsList.html I think they have some decent practice problems for integration also. What else is good? Uhhh KhanAcademy, tons of stuff on youtube. I found MIT's Calculus series really helpful. Blah I dunno >.<

OpenStudy (katielong):

okay thank you, you've been really helpful!